GIVEN: INSIDE VARIABLE COST/UNIT INSIDE FIXED COSTS 28 265,000 OUTSIDE VARIABLE COST/UNIT OUTSIDE FIXED COSTS 37 224,500 REQUcan someone help explain the proof of this problem?
Answer #1

Break even units = Difference in Fixed costs/Difference in Variable cost per unit

= (265000-224500)/(37-28)

= 4,500 units

Proof is as follows:

Inside cost = 265000 + 28*4500 = $391000

Outside cost = 224,500 + 37*4500 = $391,000

At 2500 units, inside cost = 265000 + 28*2500 = $335,000

Outside = 224500 + 37*2500 = $317,000

i.e. difference between costs = $18000

Units have been calculated as (265000-224500-18000)/(37-28)

= 2,500 units